OpenStudy (anonymous):

solve the inequality -12>=24x

OpenStudy (anonymous):

\[ -12 \ge 24x \quad \implies \quad 24x \le -12 \] Divide both sides by 24 \[ x \le \frac{-12}{24} \quad \implies \quad x \le - \frac{1}{2} \]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

sure np